Determine whether there are three consecutive integers such that their sum is three times the second integer.
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Determine whether there are three consecutive integers such that their sum is three times the second integer.
-------------
Use n-1, n & n+1.
What do you think?

Answer by ikleyn(53763)

.

For ANY three consecutive integers, their sum is three times the second number.


    Indeed, let  (n-1), n and (n+1) are three consecutive integers.


    Add them


        (n-1) + n + (n+1) = n-1 + n + n+1 = n + n + n = 3n,


    and you will get exactly what I stated.